# A Physical Model-Guided Framework for Underwater Image Enhancement and Depth Estimation

### v1

@@ III-B Deep Degradation Model
Figure 4: Detailed structure of the three sub-networks in the deep degradation model.
@@
Most previous works assume that the veiling light $B^{\infty}$ can be represented by a monochrome image with identical pixel values across spatial locations. However, this assumption fails in deep-sea scenarios where the light primarily comes from artificial point light sources. As shown in the two examples on the far right of Fig. 1, deep-sea images suffer from non-uniform illumination phenomena. Therefore, it is necessary to design a more reasonable veiling light estimation method to adapt to various lighting conditions from shallow to deep-sea environments. Fourmer [52] demonstrated that brightness, as a global feature, is primarily preserved in the center of the amplitude component of the image. Inspired by this idea, we first perform a Discrete Fourier Transform (DFT) on the underwater image to obtain the amplitude and phase components. Then we obtain the initial veiling light estimation by performing an Inverse DFT (IDFT) on the phase component and the amplitude component after applying a low-pass filter (LPF). To introduce flexibility and adjustability, the initial estimation is fed into convolutional modules with a residual connection to obtain the final veiling light $B^{\infty}$. Fig 5 compares our method with IBLA [32] and GDCP [51], both of which represent veiling light as monochrome images. For the first underwater image, the estimated veiling light maps of the three methods are similar. However, for non-uniform illumination images in deep-sea scenarios, the other two methods fail while our method achieves relatively accurate veiling light estimation.
@@
Scene depth is crucial for the physical imaging model, but accurately estimating it from a single underwater image is very challenging. A recent work, Depth Anything [45], demonstrates strong performance across various complex scenarios, including underwater environments, due to its pre-training on large-scale datasets. Depth Anything includes a DINOv2 encoder [53] to extract image features and a dense prediction Transformer (DPT) decoder [54] to predict relative depth, i.e., inverse depth. We aim to transfer the powerful zero-shot depth estimation capability of Depth Anything to our depth estimation sub-network (DEN). Specifically, we adopt the same encoder-decoder architecture and set the Dinov2 encoder to the small version of ViT [55] for efficiency. We initialize DEN with the weights from Depth Anything. During training, we freeze the DINOv2 encoder weights and only fine-tune the DPT decoder. Ultimately, DEN can accurately estimate the normalized inverse depth map $z$.
Figure 7: The structural diagram of UIEConv. UIEConv includes the local branch on the left and the global branch on the right. We annotate the shapes of the intermediate features, where $C$, $H$, and $W$ represent the number of channels, height, and width, respectively. $k$ and $s$ respectively represent the kernel size and stride in Conv2d.
@@
DEN estimates a relative inverse depth map, but what we need is the absolute depth map. To obtain the absolute depth, previous methods used predefined maximum and minimum depth values, such as 0.1-6m, to scale the relative depth of all images to the same range. However, the depth ranges of different underwater images can vary significantly, as illustrated in Fig. 6. Therefore, we introduce a factors estimation sub-network (FEN) to estimate the scaling factor $z^{Scale}$ and shift factor $z^{Shift}$ for the relative inverse depth. The absolute depth $d$ is then computed as $d=\frac{1}{z\cdot z^{Scale}+z^{Shift}}$. Additionally, FEN is responsible for estimating the scattering and attenuation coefficients. Specifically, the underwater image $I$, the estimated veiling light map $B^{\infty}$, and the inverse depth map $z$ are concatenated along the channel dimension and then input into FEN, which consists of a backbone and four structurally identical yet independent heads. The backbone includes four simple convolutional layers, each comprising convolution, instance normalization, and a ReLU activation function. The four independent heads are designed to estimate $z^{Scale}$, $z^{Shift}$, $e^{-\beta^{D}}$, and $e^{-\beta^{B}}$, respectively. Given that the attenuation coefficient $\beta^{D}$ and scattering coefficient $\beta^{B}$ are positive values, a Sigmoid activation function is applied to produce values between 0 and 1 as the terms $e^{-\beta^{D}}$ and $e^{-\beta^{B}}$. To more reasonably scale the relative depth, we add 0.1 to the Sigmoid output to generate values greater than 0.1 as $z^{Shift}$ and multiply the Sigmoid output by 2 to generate values between 0 and 2 as $z^{Scale}$. The approach allows the maximum and minimum absolute depths to vary significantly within a range of 0-10m.
@@ III-C UIEConv Model
Although the enhancement model $\mathcal{F}$ in our framework can be any advanced UIE model, such as UShape [7] or UIEC2Net [11], we have designed a simple yet more effective fully convolutional model called UIEConv. As shown in Fig. 7, UIEConv includes two branches: a global branch and a local branch.
@@
The global branch follows the common U-Net [56] encoder-decoder architecture. In the encoding part, downsampling reduces the image resolution, while in the decoding part, upsampling gradually restores the image to its original resolution. We use convolutions with a stride and kernel size of 2 for downsampling and bilinear interpolation for upsampling. The core module of the global branch is the Modern Convolutional Block (MCB) derived from the advanced vision convolutional backbone ConvNeXt [57], which employs large kernel convolutions (kernel size of 7) and a transformer-like [55] design. We replace Layer Normalization [58] with Instance Normalization [59], which has been proven to be more suitable for image restoration tasks [60]. The combination of the U-Net-like architecture and large kernel convolutions results in a very large receptive field, allowing the capture of global features in the image without the need for complex attention mechanisms.
@@ III-D Loss Function
Depth Consistency Loss
$\mathcal{L}_{depth}=\frac{1}{H*W}\sum_{i}^{H}\sum_{j}^{W}\rho(z_{ij},\hat{z}_{ij}),$ (5)
The overall loss function is a weighted sum of the above three losses:
where $\lambda_{1}$ and $\lambda_{2}$ are 0.2 and 1, respectively.
IV Experiments
IV-A Datasets and Evaluation Metrics
The shallow water datasets include LSUI [7], UIEB [5] and U45 [61]. LSUI contains 4279 pairs of real-world underwater images and clear reference images. We use 3879 pairs of these images as the training set, and the remaining 400 pairs of images as the test set. To evaluate generalization, we randomly select 90 images with reference images in UIEB as the second test set, named T90. In addition, UIEB also includes a set of 60 challenge images (C60) that do not have corresponding reference images. And U45 contains 45 carefully selected underwater images, serving as an important benchmark for UIE. For all the supervised methods, we train the models on the training set of LSUI and test them on the LUSI test set, T90, C60, and U45. The deep-sea datasets are UIID [40] and OceanDark [62]. UIID contains 3486 pairs of non-uniform illumination images with reference images. This dataset consists primarily of synthetic images, with a small portion of real images. We randomly selecte 3136 pairs for the training set and reserve the remaining 350 pairs for the test set. OceanDark comprises 183 underwater images captured by video cameras located in profound depths using artificial lighting. For all supervised methods, we train the models on the UIID training set and evaluate them on both the UIID test set and OceanDark. To quantitatively evaluate depth estimation, we conduct evaluations using the Sea-thru’s D3 and D5 subsets [35], which includes underwater images and corresponding depth maps obtained via the Structure-from-Motion algorithm.
IV-B Implementation Details
TABLE I: Quantitative comparison of various UIE methods across four datasets primarily consisting of shallow-water images. The best and second-best results are highlighted in red and blue, respectively.
@@ IV-B Implementation Details
Figure 8: Enhancement results of five images from the LSUI and UIEB dataset. The first three images come from LSUI, while the last two come from UIEB. We enlarge the local areas of the enhanced images to compare the details.
@@ IV-C Performance Evaluation
We use the training set of LSUI for training and the LSUI test set, T90, C60, and U45 for testing. Almost all the images in these datasets come from shallow-sea scenes, where the light mainly come from the sun. In Table I, we report not only the results of training UIEConv integrated into our framework (last row) but also the results of training UIEConv alone (second to last row). For datasets with reference images like LSUI and T90, our method achieves the best PSNR and SSIM metrics. For images without reference images in C60 and U45, Fusion achieves the best UCIQE, while FUnIEGAN and UIEC2Net perform best on UIQM. However, as shown in Figs. 8910, their enhancement effects are not satisfactory from a human visual perspective. This observation suggests that these two non-reference quality assessment metrics sometimes do not align with human visual perception, as noted in many previous works [5, 19, 60]. In contrast, our model consistently produces enhanced images with no color cast and clear details across various underwater conditions and degradation effects. In the last two examples, our enhanced images even appear more satisfactory than the reference images, removing excessive blue and yellow tones. Additionally, the UIEConv model has a small number of parameters and runs quickly. While integrating UIEConv into our framework reduces efficiency, it significantly improves performance. Future work will focus on designing more efficient sub-network structures.
Figure 9: Enhancement results of two images from the C60 dataset. The intensity distribution of the RGB channels is presented in boxplots. The black points represent outliers in each channel, and the gray rectangles indicate the areas where outliers are clustered.
Challenging Scene
Severe Color Distortion Scene
TABLE II: Quantitative comparison of various UIE methods across two datasets featuring deep-sea images with uneven lighting. The best and second-best results are highlighted in red and blue, respectively.
@@
For deep-sea scenarios, we train on the UIID training set and test on the UIID test set and OceanDark dataset. Table II compares our method with several UIE methods tailored for deep-sea environments with uneven lighting conditions. Our method achieves superior performance on two full-reference metrics, PSNR and SSIM. While some methods outperform ours on non-reference metrics, a visual comparison in Fig. 11 reveals the opposite. For instance, ICSP excels in the UCIQE metric but produces overexposed images. In contrast, our method enhances brightness in dark regions, mitigating uneven lighting without introducing distortion. This highlights the challenge of accurately evaluating deep-sea image quality using existing non-reference metrics. Developing appropriate non-reference metrics for deep-sea images remains an important research direction.
@@
We visualize inverse depth maps estimated by several methods in Fig. 12. Methods like GDCP and IBLA, which use visual priors to estimate transmission and depth maps, perform the worst. PUGAN struggles in texture-rich scenes. Depth Anything exhibits strong generalization ability and adapts well to underwater scenes, but its depth maps are overly smooth. In comparison, our depth estimation sub-network produces more detailed depth maps, such as the branches in the first example and the sea urchins in the second example. Additionally, we quantitatively compare the performance of Depth Anything and our method on a real underwater image dataset with depth annotations in Table III. Our method outperforms Depth Anything. Although our depth estimation sub-network is initialized from Depth Anything, through joint training with underwater image enhancement and physical imaging model parameter estimation, our method is better suited for underwater images.
TABLE III: Quantitative comparisons of depth estimation on real underwater images from two subsets of Sea-thru dataset. DA is an abbreviation for Depth Anything.
@@ IV-D Ablation Studies
To further demonstrate the effectiveness of our framework, we use it to train other advanced UIE models. As shown in Table IV, we train UIEC2Net and Ushape on the LSUI dataset, and NUICNet and IACC on the UIID dataset. We find that integrating these models into our framework yields significantly superior results compared to training them alone. For instance, PSNR increases by up to 2.4481 and SSIM increases by up to 0.0772. These experimental results indicate that our framework can be integrated with any advanced UIE model, consistently achieving performance enhancements.
TABLE IV: Comparison of several advanced UIE models before and after integration into our framework. Red numbers indicate the performance improvements achieved by our framework.
@@
TABLE V: Ablation study of the UIEConv model structure.
LSUI T90
@@
TABLE VI: Ablation study of the depth degradation model (DDM) structure and loss function on the LUSI dataset.
@@
We perform an ablation study on the structure of UIEConv, as presented in Table V. The performance significantly declines when either the global branch or the local branch of UIEConv is removed. This highlights the importance of both global features and local features for image enhancement. Consequently, the dual-branch structure of UIEConv is key to its superiority over other advanced UIE models.
@@
We conduct a comprehensive ablation study on the design details of each sub-network within the DDM, as presented in Table VI. ”Full Framework” represents the complete version of our proposed framework. We remove each unique design element in the three sub-networks one by one to observe performance changes. For VLEN, we remove its low-pass filter (w/o lowpass) or the subsequent convolutional transformation (w/o transform). For DEN, we explore three different settings during training: fine-tuning its encoder and decoder (finetune all), freezing its encoder and decoder (freeze all), and training it from scratch without initializing with Depth Anything weights (scratch). For FEN, we experiment with several different structures. First, we retain one head to simultaneously estimate the attenuation coefficient $\beta^{D}$ and the scattering coefficient $\beta^{B}$ ($\beta^{D}=\beta^{B}$). Second, we remove the heads that estimate the scale $z^{Scale}$ and shift $z^{Shift}$ of the relative depth (w/o $z^{Scale},z^{Shift}$), and instead use a predefined depth range (0.1m-10m) to scale the relative depth. Finally, we exclude the estimated veiling light $B^{\infty}$ and inverse depth $z$ as inputs to FEN (w/o additional inputs), using only the underwater image as the input. We find that removing any module or changing the training settings in the sub-networks leads to less accurate estimation performance, thereby reducing enhancement performance. Notably, training DEN from scratch causes the training process to fail to converge (resulting in NaNs), which indicates that the knowledge transferred from Pre-trained Depth Anything is crucial for the training of DEN and the entire framework. Additionally, we study the role of two additional loss functions. Training without the physical constraint loss $\mathcal{L}_{phy}$ (w/o $\mathcal{L}_{phy}$) significantly decreased performance, highlighting the importance of physical guidance in our framework. Training without the depth consistency loss $\mathcal{L}_{depth}$ (w/o $\mathcal{L}_{depth}$) also slightly reduced performance.

### v2

@@ III-B Deep Degradation Model
Figure 4: Detailed structure of the three sub-networks in the deep degradation model, where $B^{\infty}$ represents the veiling light map, $e^{-\beta^{D}}$ and $e^{-\beta^{B}}$ denote the attenuation and scattering coefficients, $z$ indicates the disparity, while $z^{Scale}$ and $z^{Shift}$ are the scale and shift factors for depth transformation. The metric depth $d$ is calculated from $z$, $z^{Scale}$ and $z^{Shift}$.
@@
Most previous works assume that the veiling light $B^{\infty}$ can be represented by a uniform color image with identical pixel values across spatial locations. However, this assumption fails in deep-sea scenarios where the light primarily comes from artificial point light sources. As shown in the two examples on the far right of Fig. 1, deep-sea images suffer from non-uniform illumination phenomena. Therefore, it is necessary to design a more reasonable veiling light estimation method to adapt to various lighting conditions from shallow to deep-sea environments. Fourmer [51] demonstrated that brightness, as a global feature, is primarily preserved in the center of the amplitude component of the image. Inspired by this idea, we first perform a Discrete Fourier Transform (DFT) on the underwater image to obtain the amplitude and phase components. Then we obtain the initial veiling light estimation by performing an Inverse DFT (IDFT) on the phase component and the amplitude component after applying a low-pass filter (LPF). To introduce flexibility and adjustability, the initial estimation is fed into convolutional modules with a residual connection to obtain the final veiling light $B^{\infty}$. Fig. 5 compares our method with IBLA [30] and GDCP [50], both of which represent veiling light as uniform color images. For the first underwater image, the estimated veiling light maps of the three methods are similar. However, for the non-uniform illumination image in deep-sea scenarios, the other two methods fail while our method achieves relatively accurate veiling light estimation.
@@
Figure 6: The structural diagram of UIEConv. UIEConv includes the local branch on the left and the global branch on the right. We annotate the shapes of the intermediate features, where $C$, $H$, and $W$ represent the number of channels, height, and width. $k$ and $s$ represent the kernel size and stride in Conv2d.
@@
Scene depth is crucial for the physical imaging model, but accurately estimating it from a single underwater image is very challenging. A recent work, Depth Anything [44], demonstrates strong performance across various complex scenarios, including underwater environments, due to its pre-training on large-scale datasets. Depth Anything includes a DINOv2 encoder [52] to extract image features and a dense prediction Transformer (DPT) decoder [53] to predict disparity, i.e., relative inverse depth. We aim to transfer the powerful zero-shot depth estimation capability of Depth Anything to our depth estimation sub-network (DEN). Specifically, we adopt the same encoder-decoder architecture and set the Dinov2 encoder to the small version of ViT [54] for efficiency. We initialize DEN with the weights from Depth Anything. During training, we freeze the DINOv2 encoder weights and only fine-tune the DPT decoder. Ultimately, DEN can accurately estimate the normalized inverse depth map $z$.
@@
DEN estimates a relative inverse depth map, but what we need is the metric (absolute) depth map. To obtain the absolute depth $d$, previous methods used predefined maximum and minimum depth values, such as 0.1-6m, to scale the relative depth of all images to the same range. However, the depth ranges of different underwater images can vary significantly, as illustrated in Fig. 7. Therefore, we introduce a factors estimation sub-network (FEN) to estimate the scaling factor $z^{Scale}$ and shift factor $z^{Shift}$ for the relative inverse depth. The absolute depth $d$ is then computed as $d=\frac{1}{z\cdot z^{Scale}+z^{Shift}}$. Additionally, FEN is responsible for estimating the scattering and attenuation coefficients. Specifically, the underwater image $I$, the estimated veiling light map $B^{\infty}$, and the inverse depth map $z$ are concatenated along the channel dimension and then input into FEN, which consists of a backbone and four structurally identical yet independent heads. The backbone extracts general image degradation features, while the four independent heads specialize in mapping these general features to different parameters to be estimated, thereby achieving a decoupled estimation. Specifically, the backbone includes four simple convolutional layers, each comprising convolution, instance normalization, and a ReLU activation function. The four independent heads are designed to estimate $z^{Scale}$, $z^{Shift}$, $e^{-\beta^{D}}$, and $e^{-\beta^{B}}$, respectively. Given that the attenuation coefficient $\beta^{D}$ and scattering coefficient $\beta^{B}$ are positive values, a Sigmoid activation function is applied to produce values between 0 and 1 as the terms $e^{-\beta^{D}}$ and $e^{-\beta^{B}}$. To more reasonably scale the relative depth, we add 0.1 to the Sigmoid output to generate values greater than 0.1 as $z^{Shift}$ and multiply the Sigmoid output by 2 to generate values between 0 and 2 as $z^{Scale}$. The approach allows the maximum and minimum absolute depths to vary significantly within a range of 0-10m.
@@ III-C UIEConv Model
Although the enhancement model $\mathcal{F}$ in our framework can be any advanced UIE model, such as UShape [7] or UIEC2Net [11], we have designed a simple yet more effective fully convolutional model called UIEConv. As shown in Fig. 6, UIEConv includes two branches: a global and a local branch.
@@
The global branch follows the common U-Netencoder-decoder architecture. In the encoding part, downsampling reduces the image resolution, while in the decoding part, upsampling gradually restores the image to its original resolution. We use convolutions with a stride and kernel size of 2 for downsampling and bilinear interpolation for upsampling. The core module of the global branch is the Modern Convolutional Block (MCB) derived from the advanced vision convolutional backbone ConvNeXt [55], which employs large kernel convolutions (kernel size of 7) and a transformer-like [54] design. We replace Layer Normalization [56] with Instance Normalization [57], which has been proven to be more suitable for image restoration tasks [58]. The combination of the U-Net-like architecture and large kernel convolutions results in a very large receptive field, allowing the capture of global features in the image without the need for complex attention mechanisms.
@@ III-D Loss Function
TABLE II: Quantitative comparison of various UIE methods across four datasets primarily consisting of shallow-water images. The best and second-best results are highlighted in red and blue, respectively.
@@
Figure 8: Enhancement results of five images from the LSUI and UIEB dataset. The first three images come from LSUI, while the last two come from UIEB.
Depth Consistency Loss
$\mathcal{L}_{depth}=\frac{1}{HW}\sum_{i}^{H}\sum_{j}^{W}\rho(z_{ij},\hat{z}_{ij}),$ (5)
The overall loss function is a weighted sum of the above three losses:
where $\lambda_{1}$ and $\lambda_{2}$ are empirically set to 0.2 and 1, respectively.
IV Experiments
IV-A Datasets and Evaluation Metrics
The shallow water datasets include LSUI [7], UIEB [5] and U45 [60]. LSUI contains 4279 pairs of real-world underwater images and clear reference images. Following [7], we use 3879 pairs of these images as the training set, and the remaining 400 pairs of images as the test set. To evaluate generalization, we use 890 images with corresponding high-quality reference images in UIEB as the second test set. In addition, UIEB also includes a set of 60 challenge images (C60) that do not have corresponding reference images. And U45 contains 45 carefully selected underwater images, serving as an important benchmark for UIE. For all the supervised methods, we train the models on the training set of LSUI and test them on the LUSI test set, UIEB, C60, and U45. The deep-sea datasets are UIID [39] and OceanDark [61]. UIID contains 3486 pairs of non-uniform illumination images with reference images. This dataset consists primarily of synthetic images, with a small portion of real images. We randomly selecte 3136 pairs for the training set and reserve the remaining 350 pairs for the test set. OceanDark comprises 183 underwater images captured by video cameras located in profound depths using artificial lighting. For all supervised methods, we train the models on the UIID training set and evaluate them on both the UIID test set and OceanDark. To quantitatively evaluate depth estimation, we conduct evaluations using the Sea-thru’s D3 and D5 subsets [33], which includes underwater images and corresponding depth maps obtained via the Structure-from-Motion algorithm.
IV-B Implementation Details
@@ IV-C Performance Evaluation
Figure 9: Enhancement results of two images from the C60 dataset. The intensity distribution of the RGB channels is presented in horizontally placed box plots which intuitively display the upper edge, upper quartile, median, lower quartile, and lower edge of each channel. The black points represent outliers in each channel, and the gray rectangles indicate the areas where outliers are clustered.
@@
We use the training set of LSUI for training and the LSUI test set, UIEB, C60, and U45 for testing. Almost all the images in these datasets come from shallow-sea scenes, where the light mainly come from the sun. In Table II, we report not only the results of training UIEConv integrated into our framework (last row) but also the results of training UIEConv alone (second to last row). For datasets with reference images like LSUI and UIEB, our method achieves the best PSNR and SSIM metrics. For images without reference images in C60 and U45, Fusion achieves the best UCIQE, while FUnIEGAN and UIEC2Net perform best on UIQM. However, as shown in Figs. 8910, their enhancement effects are not satisfactory from a human visual perspective. This observation suggests that these two non-reference quality assessment metrics sometimes do not align with human visual perception, as noted in many previous works [5, 16, 58]. In contrast, our model consistently produces enhanced images with no color cast and clear details across various underwater conditions and degradation effects. In the last two examples, our enhanced images even appear more satisfactory than the reference images, removing excessive blue and yellow tones. Additionally, the UIEConv model has a small number of parameters and runs quickly. While integrating UIEConv into our framework reduces efficiency, it significantly improves performance. Future work will focus on designing more efficient sub-network structures.
TABLE III: Quantitative comparison of various UIE methods across two datasets featuring deep-sea images with uneven lighting. The best and second-best results are highlighted in red and blue, respectively.
@@
Challenging Scene
Severe Color Distortion Scene
@@
For deep-sea scenarios, we train on the UIID training set and test on the UIID test set and OceanDark dataset. Table III compares our method with several UIE methods tailored for deep-sea environments with uneven lighting conditions. Our method achieves superior performance on two full-reference metrics, PSNR and SSIM. While some methods outperform ours on non-reference metrics, a visual comparison in Fig. 11 reveals the opposite. For instance, ICSP excels in the UCIQE metric but produces overexposed images. In contrast, our method enhances brightness in dark regions, mitigating uneven lighting without introducing distortion. This highlights the challenge of accurately evaluating deep-sea image quality using existing non-reference metrics. Developing appropriate non-reference metrics for deep-sea images remains an important research direction.
@@
We visualize inverse depth maps estimated by several methods in Fig. 12. Methods like GDCP and IBLA, which use visual priors to estimate transmission and depth maps, perform the worst. PUGAN struggles in texture-rich scenes. Depth Anything exhibits strong generalization ability and adapts well to underwater scenes, but its depth maps are overly smooth. In comparison, our depth estimation sub-network produces more detailed depth maps, such as the branches in the first example and the sea urchins in the second example. Additionally, we quantitatively compare the performance of Depth Anything and our method on a real underwater image dataset with depth annotations in Table IV. Our method outperforms Depth Anything. Although our depth estimation sub-network is initialized from Depth Anything, through joint training with underwater image enhancement and physical imaging model parameter estimation, our method is better suited for underwater images.
TABLE IV: Quantitative comparisons of depth estimation on real underwater images from two subsets of Sea-thru dataset. DA is an abbreviation for Depth Anything.
@@ IV-D Ablation Studies
To further demonstrate the effectiveness of our framework, we use it to train other advanced UIE models. As shown in Table V, we train UIEC2Net and Ushape on the LSUI dataset, and NUICNet and IACC on the UIID dataset. To ensure a fair comparison, we keep all the model and training hyperparameters unchanged when integrating the model into our framework. We find that integrating these models into our framework yields significantly superior results compared to training them alone. For instance, PSNR increases by up to 2.4481 and SSIM increases by up to 0.0772. These experimental results indicate that our framework can be integrated with any advanced UIE model, consistently achieving performance enhancements.
TABLE V: Comparison of several advanced UIE models before and after integration into our framework. Red numbers indicate the performance improvements achieved by our framework.
@@
TABLE VI: Ablation study of the UIEConv model structure.
LSUI UIEB
@@
TABLE VII: Ablation study of the depth degradation model (DDM) structure and loss function on the LUSI dataset.
@@
We perform an ablation study on the structure of UIEConv, as presented in Table VI. The performance significantly declines when either the global branch or the local branch of UIEConv is removed. This highlights the importance of both global features and local features for image enhancement. Consequently, the dual-branch structure of UIEConv is key to its superiority over other advanced UIE models.
@@
We conduct a comprehensive ablation study on the design details of each sub-network within the DDM, as presented in Table VII. ”Full Framework” represents the complete version of our proposed framework. We remove each unique design element in the three sub-networks one by one to observe performance changes. For VLEN, we remove its low-pass filter (w/o lowpass) or the subsequent convolutional transformation (w/o transform). For DEN, we explore three different settings during training: fine-tuning its encoder and decoder (finetune all), freezing its encoder and decoder (freeze all), and training it from scratch without initializing with Depth Anything weights (scratch). For FEN, we experiment with several different structures. First, we retain one head to simultaneously estimate the attenuation coefficient $\beta^{D}$ and the scattering coefficient $\beta^{B}$ ($\beta^{D}=\beta^{B}$). Second, we remove the heads that estimate the scale $z^{Scale}$ and shift $z^{Shift}$ of the relative depth (w/o $z^{Scale},z^{Shift}$), and instead use a predefined depth range (0.1m-10m) to scale the relative depth. Finally, we exclude the estimated veiling light $B^{\infty}$ and inverse depth $z$ as inputs to FEN (w/o additional inputs), using only the underwater image as the input. We find that removing any module or changing the training settings in the sub-networks leads to less accurate estimation performance, thereby reducing enhancement performance. Notably, training DEN from scratch causes the training process to fail to converge (resulting in NaNs), which indicates that the knowledge transferred from Pre-trained Depth Anything is crucial for the training of DEN and the entire framework. Additionally, we study the role of two additional loss functions. Training without the physical constraint loss $\mathcal{L}_{phy}$ (w/o $\mathcal{L}_{phy}$) significantly decreased performance, highlighting the importance of physical guidance in our framework. Training without the depth consistency loss $\mathcal{L}_{depth}$ (w/o $\mathcal{L}_{depth}$) also slightly reduced performance.